OpenStudy (loser66):

Help me out please. \(x'+\dfrac{2}{(150+0.5t)}x =7.5\) I got trouble with \(e^{\int p(t) dt}\)

OpenStudy (loser66):

\(\int p(t) dt = \int \dfrac{2}{150+0.5t}dt\) Case 1: \(dfrac{2}{150+0.5t}=\dfrac{4}{300+t}\) Hence when taking integral, I got \(4 ln(300+t)\), when plugging in, \(\mu= (300+t)^4\)

OpenStudy (loser66):

Case2, If I let it as it is, Let u = (150+0.5t) then 2du = dt therefore \(\int \dfrac{4du}{u}=4ln u= 4ln(150+0.5t)\) Plugging in, I got \(\mu= (150+0.5t)^4\) OHOHOH>>>> NO. What is wrong?

OpenStudy (mww):

hint: everyone forgets about the constant of integration... Is not ln(ax) = ln(x)+ln(a)?

OpenStudy (loser66):

I don't get what you mean. We don't have ln (ax), we have a*ln x

OpenStudy (mww):

ok try this problem. Is integral of ln(x) then same as ln(ax) for some constant a?

OpenStudy (mww):

\[\int\limits \ln x dx = \frac{ 1 }{ x } + C \] \[\int\limits \ln(ax)dx = a(\frac{ 1 }{ ax }) +C = \frac{ 1 }{ x } + C\] So there's no difference really)

OpenStudy (loser66):

@ganeshie8 help me please.

ganeshie8 (ganeshie8):

Yeah there can be infinitely many integrating factors, just need one to use the product rule in reverse.

OpenStudy (loser66):

no, ganeshie8, the result must be the same so that we can get the correct solution.

ganeshie8 (ganeshie8):

All give you the correct solution

OpenStudy (loser66):

\((300+t)^4\neq (150+0.5t)^4\)

ganeshie8 (ganeshie8):

Who said they need to be equal ?

OpenStudy (loser66):

At the end, we put initial condition to solve for C.

ganeshie8 (ganeshie8):

Right, all integrating factors give you the same solution for IVP

OpenStudy (loser66):

And if they are different, I have 2 different C's . How can I know which one is correct one?

ganeshie8 (ganeshie8):

The particular value of C will change based on the general solution

OpenStudy (loser66):

IC x(0) =20 hence , the first one gives me C =-430 (300)^4 the second one gives me C =-430(150)^4 OHOH, no

ganeshie8 (ganeshie8):

And what are the general solutions ?

OpenStudy (loser66):

For the first one: \(x(t)=1.5(300+t)+\dfrac{C}{(300+t)^4}\)

OpenStudy (loser66):

The second one is \(x(t) =3(150+0.5t)+\dfrac{C}{(150+0.5t)^4}\)

OpenStudy (mww):

Yeah the C can change so don't think C1 = C2. common mistake.

ganeshie8 (ganeshie8):

Is it hard to see that you get identical equations after you plugin the respective C's ?

OpenStudy (loser66):

Oh, I got you. MY GOD, how stupid I am. Thanks a ton.

OpenStudy (loser66):

yes, yes, yes, lalala... hahaha... I am struggling with it more than 2 hours.

ganeshie8 (ganeshie8):

Haha nope. you're not stupid. lookup existence and uniqueness theorem. It's proof is hard but you might enjoy :)

OpenStudy (loser66):

I have final on next Monday, must be ok with all of the materials.

ganeshie8 (ganeshie8):

good luck with the exam !

OpenStudy (loser66):

Thank you :)